OpenStudy (anonymous):

how do you solve the equation: f=12gh solve for h

OpenStudy (saifoo.khan):

\[h = \frac{f}{12g}\]

OpenStudy (agreene):

f=12*g*h divide off your 12 and g f/12g = h

OpenStudy (anonymous):

thanks!!!!

OpenStudy (anonymous):

P=n+k+m solve for n

OpenStudy (agreene):

P=N+K+M subtract the stuff across the = sign you dont want (K,M) and get P-K-M=N